OpenStudy (anonymous):

Write the ordered pair solution like this (1,2) with no spaces, or write no solution or infinitely many solutions. x + y = -8 and x - y = 8

OpenStudy (anonymous):

solution is (0,-8)

OpenStudy (anonymous):

Uuughh. i still dont get it.

OpenStudy (anonymous):

adding both equations we get 2x=0 this emplies x=0 put it in equation 2 or 1 you will get -8

OpenStudy (anonymous):

its too simple just add two equations

OpenStudy (anonymous):

wait did u use (1,2) too get the answer.

OpenStudy (anonymous):

\[x+y=-8\] + \[x-y=8\] gives 2x=0 divide by 2 on both sides you get x=0 put this in \[x+y=-8\] u get 0+y=-8 or y=-8 then ans will be (x,y)=(0,-8)

OpenStudy (anonymous):

(1,2) is not its solution

OpenStudy (anonymous):

Ok thanksss : )